# E-values, Multiple Testing and Beyond

### v1

@@ 4.1.1 Choice of significance levels and weights
We now discuss the choices of the significance levels $(\alpha_{\mathrm{BH}},\alpha_{\mathrm{BC}})$ and the weights $(w_{\mathrm{BH},i},w_{\mathrm{BC},i})$ in Algorithm 1, which play important roles in the hybrid procedure. Different from the target FDR level $\alpha_{\mathrm{eBH}}$, the choice of $(\alpha_{\mathrm{BH}},\alpha_{\mathrm{BC}})$ does not affect the FDR control level but instead affect the power of the hybrid procedure. Following the discussions in Section 3.2 of Ren and Barber (2022), when there are $n_{a}$ non-nulls with extremely strong signals, we expect that
$1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\approx\sum_{j\in\mathcal{H}_{0}}\mathbbm{1}\{p_{j}\leq T_{\mathrm{BC}}\}\approx\frac{\alpha_{\mathrm{BC}}n_{a}}{1-\alpha_{\mathrm{BC}}}.$
In a similar spirit, we expect the FDR of the BH procedure to be $\tau_{0}\alpha$, where $\tau_{0}=n_{0}/n.$ Let $R_{\mathrm{BH}}$ be the number of rejections in the BH procedure. Then we have $R_{\mathrm{BH}}\approx n_{a}/(1-\tau_{0}\alpha_{\mathrm{BH}})$ and
$T_{\mathrm{BH}}=\frac{\alpha_{\mathrm{BH}}R_{\mathrm{BH}}}{n}\approx\frac{\alpha_{\mathrm{BH}}n_{a}}{n(1-\tau_{0}\alpha_{\mathrm{BH}})}.$
Thus the hybrid procedure will reject $H_{i}$ with $i\notin\mathcal{H}_{0}$ when
$e_{i}\approx\frac{\omega_{\mathrm{BC},i}n(1-\alpha_{\mathrm{BC}})}{\alpha_{\mathrm{BC}}n_{a}}+\frac{\omega_{\mathrm{BH},i}n(1-\tau_{0}\alpha_{\mathrm{BH}})}{\alpha_{\mathrm{BH}}n_{a}}\geq\frac{n}{\alpha_{\mathrm{eBH}}n_{a}},$
which is equivalent to
$\frac{\omega_{\mathrm{BC},i}(1-\alpha_{\mathrm{BC}})}{\alpha_{\mathrm{BC}}}+\frac{\omega_{\mathrm{BH},i}(1-\tau_{0}\alpha_{\mathrm{BH}})}{\alpha_{\mathrm{BH}}}\geq\frac{1}{\alpha_{\mathrm{eBH}}}.$ (13)
@@
We found that setting $\alpha_{\mathrm{BC}}=\alpha_{\mathrm{BH}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$ fulfills the above constraint, and leads to good performance in our numerical studies.
@@
Next, we discuss the selection of $(w_{\mathrm{BH},i},w_{\mathrm{BC},i})$ which balances the contribution from the two methods. A natural choice would be to set $w_{\mathrm{BH},i}=w_{\mathrm{BC},i}=0.5$ for all $i$. The theorem below shows that if the weights are independent of the p-values and satisfy $w_{\mathrm{BH},i}+\max_{j}w_{\mathrm{BC},j}\leq 1$ for all $i$, then the e-values defined in (12) satisfy Condition (3). Therefore, according to Proposition 1, the corresponding e-BH procedure controls the FDR at the desired level.
Theorem 4.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and super-uniform, and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. Assume that $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ are independent of the p-values, and $w_{\mathrm{BH},i}+\max_{j}w_{\mathrm{BC},j}\leq 1$ for all $i$. Then, the weighted e-values given in (12) satisfy Condition (3).
[…]
@@
As a consequence of Theorem 5, the hybrid procedure with the data-dependent weights in (14) provides finite sample FDR control. Furthermore, in view of the proof of Theorem 5, the above conclusion remains true if we replace $T_{\mathrm{BH},i}$ in (14) by any function of $\tilde{p}_{-i}$.
@@ 4.1.2 Numerical studies
We investigate the finite sample performance of the hybrid procedure via several simulation examples. We fix the sample size $n=1,000$ and set the significance level at $\alpha_{\mathrm{eBH}}=0.05$. For each experimental setting, the average FDP (which estimates the FDR) and the average power based on 500 Monte Carlo replications are reported. We consider two different ways to implement the hybrid procedure: (1) eBH_Ada, for which the weights are calculated via (14) and $\alpha_{\mathrm{BH}}=\alpha_{\mathrm{BC}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$; (2) eBH_Ave, for which the weights are set as $w_{\mathrm{BH},i}=w_{\mathrm{BC},i}=0.5$ for all $i=1,\cdots,n$ and $\alpha_{\mathrm{BH}}=\alpha_{\mathrm{BC}}=\alpha_{\mathrm{eBH}}/2$. Notice that the weights defined in (14) involve the term $T_{\mathrm{BC},j,i}$, which can be computationally expensive. To reduce the computational burden, we also consider a fast implementation of eBH_Ada, referred to as fast_eBH_Ada, which uses the weights
@@
but otherwise is the same as eBH_Ada.
We consider two settings. In Setting S1, we set $n_{a}$ to be $100$ and $200$, corresponding to $10\%$ and $20\%$ signals. For the null hypotheses, the p-values are generated independently from the uniform distribution on $[0,1]$; for the alternative hypotheses, the p-values are generated independently from the uniform distribution on $[0,p_{1}]$, where $p_{1}$ takes values between $1.25\times 10^{-2}$ and $1.5\times 10^{-2}$. Here, a larger $p_{1}$ indicates a weaker signal strength on average. Figure 1 summarizes the results for Setting S1. All methods under consideration control the FDR at the 5% level. For the dense signal case (i.e., 20% of the hypotheses are non-nulls), the BH procedure is quite conservative, while the BC procedure delivers significantly higher power. The performance of eBH_Ada is between the BH and BC procedures. In contrast, eBH_Ave does not provide much improvement over the BH procedure. It is also worth noting that fast_eBH_Ada delivers almost identical results compared to eBH_Ada but with a much lower computational cost. When $n_{a}=100$, the eBH_Ada procedure has nearly the same power as the BC procedure, and it outperforms eBH_Ave whose power is almost identical to the BH procedure.
@@
Figure 1: FDR and power for Setting S1. The left panel and the right panel correspond to the cases of $n_{a}=100$ and $n_{a}=200$, respectively. The results are based on $500$ Monte Carlo replications.
In Setting S2. we set $n_{a}$ to be $50$ and 100, corresponding to $5\%$ and $10\%$ of all hypotheses. Under the null hypothesis, we generate the test statistics $X_{i}$ from the standard normal distribution $\mathcal{N}(0,1)$. Under the alternatives, $X_{i}$ follows $\mathcal{N}(\mu\log(n),1)$, where the parameter $\mu$ controls the signal strength, which varies between $0.3$ and $0.5$. Subsequently, the p-values are computed as $p_{i}=1-\Phi(X_{i})$, where $\Phi$ denotes the distribution function of the standard normal distribution. The simulation results are presented in Figure 2. When the signal density is 5% (i.e., $n_{a}=50$), the BH procedure performs better than the BC procedure. eBH_Ada has nearly the same power as the BH procedure, and it again outperforms eBH_Ave, which slightly improves over the BC procedure. When the signal density increases to 10% (i.e., $n_{a}=100$), the performance gap between the BH and BC procedures diminishes. eBH_Ada and its fast version produce almost the same power as the BH procedure, outperforming both the eBH_Ave and BC procedures. In this case, the performance of eBH_Ave is worse than that of the BH and BC procedures.
Figure 2: FDR and power for Setting S2. The left panel and the right panel correspond to the cases of $n_{a}=50$ and $n_{a}=200$, respectively. Results are based on $500$ Monte Carlo replications.
In summary, eBH_Ada is capable of improving the worse-performing method within the BH and BC procedures under different scenarios. Interestingly, its power can be nearly identical to the better-performing method, which demonstrates the adaptivity of eBH_Ada. fast_eBH_Ada delivers almost the same results as eBH_Ada in all settings with a much lower computational burden. We therefore recommend the use of fast_eBH_Ada in practice.
4.2 Fair multiple testing
In recent years, there has been a growing interest in developing machine learning algorithms that ensure certain notion of fairness for groups categorized by protected attributes, such as gender, race, or ethnicity. However, the development of statistical inference procedures that guarantee specific notions of fairness has not received the same level of attention. To address this gap, we introduce a new notion of fairness in the context of multiple hypothesis testing, and propose a procedure to ensure this notion of fairness.
Suppose we have $n$ hypotheses which can be partitioned into $G$ distinct groups, denoted as $\{\mathcal{G}_{g}\colon g=1,\dots,G\}$, according to some protected attributes, where $\cup_{g=1}^{G}\mathcal{G}_{g}=[n]:=\{1,2,\dots,n\}$ and $\mathcal{G}_{g}\cap\mathcal{G}_{g^{\prime}}=\emptyset$ for $g\neq g^{\prime}$. Recall that $\theta_{i}\in\{0,1\}$ represents the underlying truth of $H_{i}$ and $\delta_{i}\in\{0,1\}$ denotes the decision rule for $H_{i}$. We define the group-wise FDP and FDR based on $\bm{\delta}$ respectively as:
$\mathrm{FDP}_{g}(\bm{\delta})=\frac{\sum_{i\in\mathcal{G}_{g}}(1-\theta_{i})\delta_{i}}{1\vee\sum_{i\in\mathcal{G}_{g}}\delta_{i}},\quad\text{FDR}_{g}(\bm{\delta})=\mathbb{E}[\text{FDP}_{g}(\bm{\delta})],\quad g=1,2,\dots,G.$
Definition 1.
A decision rule $\bm{\delta}$ with the target FDR level $\alpha$ is called a fair decision rule if it controls both the group-wise FDRs uniformly over $1\leq g\leq G$ and the overall FDR at the level $\alpha$, i.e., $\max_{1\leq g\leq G}\text{FDR}_{g}(\bm{\delta})\leq\alpha$ and $\text{FDR}(\bm{\delta})\leq\alpha$.
Remark 1.
Among various notions of fairness that have been discussed in the literature (Verma and Rubin, 2018), predictive parity in the classification context (Chouldechova, 2017) is most relevant to our definition of fairness. To elaborate, let us consider a binary classification problem where $Y\in\{0,1\}$ represents the true labels and $\widehat{Y}\in\{0,1\}$ denotes the predicted labels. In this scenario, the FDR is defined as $P(Y=0|\widehat{Y}=1)$. A classifier satisfies predictive parity if it ensures an equal FDR across all groups. However, the multiple testing problem differs from the classification problem as the underlying truth of each hypothesis is unobserved and cannot be used to construct the decision rule. In our context, it is more appropriate to control the FDRs across different groups at the same desired level rather than forcing them to be the same.
[…]
Algorithm A 0.039 0.006 0.067
Algorithm B 0.057 0.048 0.025
Table 1: FDR results for Setting E2, where the nominal FDR level is $5\%$.
To overcome the issues above, we consider a multiple testing procedure by assembling the e-values from the BC procedure applied to each group separately. Specifically, we implement the BC procedure at the level $\alpha$ for each individual group and let

### v2

@@ 4.1.1 Choosing Significance Levels and Weights
Different from the target FDR level $\alpha_{\mathrm{eBH}}$, the choice of $(\alpha_{\mathrm{BH}},\alpha_{\mathrm{BC}})$ does not affect the FDR control level but instead affects the power of the hybrid procedure. Following the discussion in Section 3.2 of Ren and Barber (2024), when there are $n_{a}$ non-nulls with extremely strong signals, we expect that $1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\approx\bigl{(}\alpha_{\mathrm{BC}}n_{a}\bigr{)}/\bigl{(}1-\alpha_{\mathrm{BC}}\bigr{)}$. In a similar spirit, we expect the FDR of the BH procedure to be $\tau_{0}\alpha_{\mathrm{BH}}$, where $\tau_{0}=n_{0}/n.$ Let $R_{\mathrm{BH}}$ be the number of rejections in the BH procedure. Then we have $R_{\mathrm{BH}}\approx n_{a}/(1-\tau_{0}\alpha_{\mathrm{BH}})$ and $T_{\mathrm{BH}}\approx\bigl{(}\alpha_{\mathrm{BH}}n_{a}\bigr{)}/\bigl{(}n(1-\tau_{0}\alpha_{\mathrm{BH}})\bigr{)}.$ Thus the hybrid procedure will reject $H_{i}$ with $i\notin\mathcal{H}_{0}$ when
$e_{i}\approx\frac{w_{\mathrm{BC},i}n(1-\alpha_{\mathrm{BC}})}{\alpha_{\mathrm{BC}}n_{a}}+\frac{w_{\mathrm{BH},i}n(1-\tau_{0}\alpha_{\mathrm{BH}})}{\alpha_{\mathrm{BH}}n_{a}}\geq\frac{n}{\alpha_{\mathrm{eBH}}n_{a}}.$
@@
We found that setting $\alpha_{\mathrm{BC}}=\alpha_{\mathrm{BH}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$ fulfills the above constraint if $w_{\mathrm{BC},i}+w_{\mathrm{BH},i}\leq 1$, and leads to good performance in our numerical studies.
@@
Next, we consider the selection of $(w_{\mathrm{BH},i},w_{\mathrm{BC},i})$, which balances the contributions from the two methods. A natural choice is to set $w_{\mathrm{BH},i}=w_{\mathrm{BC},i}=0.5$ for all $i$. According to Proposition 5 in Li and Zhang (2025), the e-values defined in (7) with these data-independent weights satisfy Condition (2). Therefore, by Proposition 1, the corresponding e-BH procedure controls the FDR at the desired level.
Our simulations indicate that the e-BH procedure, which is based on averaged e-values, achieves only a slight improvement in power over the less powerful method between the BH and BC procedures. To address this issue, we introduce a data-dependent approach to construct weights. Our idea is partly motivated by the leave-one-out technique used in proving the FDR control for the BH (Ferreira and Zwinderman, 2006) and BC procedures (Barber et al., 2020). For $i=1,\dots,n$, denote $\tilde{p}_{i}\coloneqq\min\{p_{i},1-p_{i}\}$, $\mathbf{p}_{-i}\coloneqq\{p_{1},\dots,p_{i-1},\,\tilde{p}_{i},\,p_{i+1},\dots,p_{n}\}$ and $\tilde{\mathbf{p}}_{-i}\coloneqq\{\tilde{p}_{1},\dots,\tilde{p}_{i-1},\,0,\,\tilde{p}_{i+1},\dots,\tilde{p}_{n}\}$. By viewing $T_{\mathrm{BH}}$ and $T_{\mathrm{BC}}$ as functions of the p-values, we define $T_{\mathrm{BH},i}=T_{\mathrm{BH}}(\tilde{\mathbf{p}}_{-i})$ and $T_{\mathrm{BC},j}=T_{\mathrm{BC}}(\mathbf{p}_{-j})$. Further define $T_{\mathrm{BC},j,i}$ in the same way as $T_{\mathrm{BC},j}$ but with $p_{i}$ being replaced by 0 when $j\neq i$. We propose the following e-value weights
@@
As a consequence of Theorem 2, the hybrid procedure with the data-dependent weights in (8) provides finite-sample FDR control. Furthermore, in view of the proof of Theorem 2, the above conclusion remains true if we replace $T_{\mathrm{BH},i}$ in (8) by any (deterministic) function of $\tilde{\mathbf{p}}_{-i}$.
@@ 4.1.2 Numerical Studies
We investigate the finite sample performance of the hybrid procedure via several simulation examples. We set the significance level at $\alpha_{\mathrm{eBH}}=0.05$. For each experimental setting, the average FDP (which estimates the FDR) and the average power based on 500 Monte Carlo replications are reported. We consider two different ways to implement the hybrid procedure: (1) eBH_Ada, for which the weights are calculated via (8) and $\alpha_{\mathrm{BH}}=\alpha_{\mathrm{BC}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$; (2) eBH_Ave, for which the weights are set as $w_{\mathrm{BH},i}=w_{\mathrm{BC},i}=0.5$ for all $i=1,\dots,n$ and $\alpha_{\mathrm{BH}}=\alpha_{\mathrm{BC}}=\alpha_{\mathrm{eBH}}/2$. Notice that the weights defined in (8) involve the term $T_{\mathrm{BC},j,i}$, which can be computationally expensive. To reduce the computational burden, we also consider a fast implementation of eBH_Ada, referred to as fast_eBH_Ada, which uses the weights
@@
but otherwise, it is the same as eBH_Ada.
We generate p-values from two settings, Setting S1 and Setting S2, where the BH procedure outperforms the BC procedure in Setting S1 while the BC procedure provides significantly higher power in Setting S2. The simulation setups are deferred to Supplement F.
The left panel of Figure 2 summarizes the results for Setting S1. All methods under consideration control the FDR at the 5% level. eBH_Ada demonstrates nearly the same power as the BH procedure and surpasses eBH_Ave, which offers a slight improvement over the BC procedure. Notably, fast_eBH_Ada achieves results almost identical to those of eBH_Ada but at a significantly lower computational cost.
The right panel of Figure 2 summarizes the results for Setting S2. The performance of eBH_Ada lies between that of the BH and BC procedures. In contrast, eBH_Ave shows little improvement over the BH procedure. Again, fast_eBH_Ada produces nearly identical results to eBH_Ada but with a much lower computational cost.
@@
Figure 2: The left panel and right panel correspond to FDR and power for Setting S1 and Setting S2, respectively. The results are based on $500$ Monte Carlo replications.
Next, we fix one setting in Setting S1 and calculate the computational cost for each method, as shown in Table F.1 in Supplement F. We observe that the data-independent method, eBH_Ave, performs the analysis as quickly as the BH or BC procedures. The fast_eBH_Ada method demonstrates acceptable speed and is significantly faster than the eBH_Ada method.
In summary, eBH_Ada effectively enhances the performance of the weaker method between the BH and BC procedures across different scenarios. Notably, its power can be nearly identical to that of the stronger of the two procedures, highlighting the adaptivity of eBH_Ada. fast_eBH_Ada achieves almost identical results to eBH_Ada in all settings, while significantly reducing computational cost. Therefore, we recommend fast_eBH_Ada for practical applications.
5 Structure-Adaptive Multiple Testing
Having access to various types of auxiliary information that reflect the structural relationships among hypotheses is becoming increasingly common. Taking advantage of such auxiliary information can improve the statistical power in multiple testing. In this section, we consider the scenarios where, in addition to the p-value $p_{i}$, there is associated structural information in the form of a covariate $x_{i}$ for each hypothesis. This side information represents heterogeneity among the p-values and may affect the prior probabilities of the null hypotheses being true or the signal strength under alternatives. Our goal is to develop a multiple testing procedure that can incorporate such external structural information to improve statistical power and guarantee FDR control in finite sample. The high-level idea behind our approach is to relax the p-value thresholds for hypotheses more likely to be non-null and tighten the thresholds for others through the use of a hypothesis-specific rejection rule, i.e., $\varphi_{i}(p_{i})\leq t$, so that the FDR can be controlled.
Our proposed method combines the cross-fitting technique (Ignatiadis and Huber, 2021) (a sample-splitting and fitting approach that enables learning the hypothesis-specific rejection function $\varphi_{i}$ without overfitting as long as the hypotheses can be partitioned into independent folds) with the FBC procedure introduced in Li and Zhang (2025). First, we randomly split the data into $G$ distinct groups, denoted as $\{\mathcal{G}_{g}\colon g=1,\dots,G\}$, where $\bigcup_{g=1}^{G}\mathcal{G}_{g}=[n]$ and $\mathcal{G}_{g}\cap\mathcal{G}_{g^{\prime}}=\emptyset$ for $g\neq g^{\prime}$. We estimate the rejection function $\varphi_{i}$ for hypothesis $i$ in the $g$th group using the data from all other groups, which ensures that the estimated rejection function is independent of the p-values in group $g$. We then apply the FBC procedure based on the estimated rejection functions separately to each group and obtain the corresponding e-values. Finally, we collect all the e-values and apply the e-BH procedure at the target level to control the FDR.
To describe the cross-fitting procedure, let us assume that $\varphi_{i}(p)=\varphi(p,x_{i};\beta)$ for some unknown parameter $\beta$ that needs to be estimated from the data. We define the cross-fitting estimate as $\hat{\beta}_{-g}=\operatorname*{arg\,min}_{\beta\in\mathcal{B}}\sum_{i\notin\mathcal{G}_{g}}\mathcal{L}(p_{i},x_{i},\beta)$. Here $\mathcal{L}$ is some loss function, such as negative log-likelihood, and $\mathcal{B}$ is a parameter space. Given $\hat{\beta}_{-g}$, we define $\hat{\varphi}_{i}(p)=\varphi(p,x_{i};\hat{\beta}_{-g})$ for $i\in\mathcal{G}_{g}$. Next, we apply the FBC procedure using the cross-fitted functions $\{\hat{\varphi}_{i}(\cdot)\}_{i\in\mathcal{G}_{g}}$ at the level $\alpha_{\mathrm{FBC}}$. The corresponding threshold for the $g$th group is given by